3. The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line? (1 point)
Answer by mamiya(56)

let's (D): Y=ax+b, (a& b, real numbers)be the equation of the line.
the x-intercept of (D)is -5, means that the point A(-5, 0) is a point of (D)
the y-intercept of (D)is -2, means that the point B(0,-2) is a point of (D)
from these data, we have, -5a+b=0
0+b=-2
b=-2
a=b/5= -2/5
so (D): y= -2/5x-2